# Non-Abelian anyons in a periodically-driven Abelian model

### v2

@@ II.2 Numerical determination of the Floquet Hamiltonian
As discussed in the main text and detailed in section ‘Vortex Exchange Phases’, the calculation of the exchange phases builds on the assumption that ground states of the effective Floquet Hamiltonian for the driven fermions, corresponding to a given distribution of $e$ quasiparticles, are prepared. In the high-frequency regime considered here, all $\hat{H}_{F,\mathrm{BdG}}^{[t_{0}]}$ for any initial time $t_{0}$ and their ground state coincide to leading order in $1/\omega$ [see Eq. (S31b)]: the Hamiltonians are both well approximated by the time-averaged Hamiltonian of Eq. (S36), which possesses entirely the desired topological properties for the effective fermion bands, and only differ by small corrections due to the weak higher-order terms. This is indeed verified by comparing the vortex exchange phases obtained using ground states of $\hat{H}_{\rm FBdG}^{[t_{0}]}$ for different initial time $t_{0}$, as shown in Fig. S4. From Fig. S4, one can appreciate that, for all system sizes shown, the variation in the phase is within $\sim 0.3\%$.
Figure S4: Vortex exchange phase $R_{\psi}^{ee}$ in the $\psi$ sector, divided by its ideal value $83\pi/$ 8 as a function of the global phase $\varphi_{0}$ of the driving functions, defined by $d_{x}(t)=J+2\Delta\cos(\omega t+\varphi_{0})$ and $d_{y}(t)=J+2\Delta\cos(\omega t+\varphi_{0}+\pi/2)$. The relative phase $\pi/2$ imprints $p+ip$ pairing, as discussed in the main text. The driving amplitude and frequency are set at $J=\Delta=\mu_{\psi}/16$ and $\omega=2\mu_{\psi}-4J=7\mu_{\psi}/4$ in the topological region. The gray dotted line is a guide for the eye at the ideal value 1.
@@ III Vortex exchange phases
with a Thouless matrix $T_{ij}$, with respect to the chosen reference state |$\ket{0_{\rm tc}}$, which is the fermion vacuum. The normalization constant is $\mathcal{N}=\sqrt{\mathrm{det}(u)}$, with $u$ defined in Eq. (S55). In case |$\ket{\Phi}$ has vanishing overlap with the fermion vacuum, $\langle\Phi|0_{\rm tc}\rangle=0$, a different reference state needs be chosen, as also further discussed below. By imposing $\hat{\alpha}_{n}\ket{\Phi}=0$, the Thouless matrix $T_{ij}$ is found as $T=-(u^{*})^{-1}v^{*}$ [53]. Since, for the toric-code model on a torus, only pairs of fermionic quasiparticles can be created or annihilated, the physical states must have even fermionic parity. The ground state parity is determined by the Floquet-BdG Hamiltonian and is computed following Ref. [8]. Namely, the Floquet Hamiltonian is first expressed in terms of Majorana operators [Eq. (S14)] as $\hat{H}_{F}=i\sum_{p,p^{\prime}}{H}^{(\gamma)}_{F,pp^{\prime}}\hat{\gamma}_{p}\hat{\gamma}_{p^{\prime}}^{\prime}$. The ground state parity is then given by the Pfaffian $\mathrm{Pf}(B)$ of the real skew-symmetric matrix $B$ defined as
@@
In case the lowest energy state of $\hat{H}_{\rm BdG}$ has odd parity, the physical ground state will be the one containing the first BdG quasiparticle. This can be determined as the state which is annihilated by all the operators $\alpha_{j}$ with $j\neq 1$ and by $\alpha_{j}^{\dagger}$ [53, 24].
Geometric phase. To compute the geometric phase accumulated by moving a vortex along a given path, we follow the strategy of Refs. [24, 50], adapted here to the Floquet system. We consider controlled vortex movement within the bulk, that never spans the whole lattice, such that no flipping of the eigenvalues of the Wilson loop operators (change of superselection sector) can occur. For each position of the vortices along the path, the fermion Floquet-BCS corresponding to that vortex configuration is computed from the BdG Hamiltonian as explained in the previous section, resulting in a sequence of fermion ground states |$\ket{\Phi_{j}}$. To avoid dependence of the phase from the details of the exchange path, Levin-Wen’s protocol [50] is employed, which requires one to compute the difference $\theta=\theta_{P}-\theta_{P^{\prime}}$ of the phases $\theta_{P}$ accumulated along two ‘T-shaped’ paths depicted in Fig. 3(a). The phase along path $P$ is computed as
@@
On the one hand, this choice of reference state ‘saves’ one computation step by fixing the phase in the first step to $\mathrm{Arg}\langle\Phi_{2}|\Phi_{1}\rangle=0$, since the ansatz of Eq. (S63) implicitly fixes the gauge along the path through the condition that $\langle\Phi_{i}|\Phi_{1}\rangle=\mathcal{N}_{1}$ is real. On the other, it is a convenient choice since $\langle\Phi_{i}|\Phi_{1}\rangle$ typically remains fairly large, avoiding numerical issues in the construction of the state associated with small overlaps.
@@
$\begin{pmatrix}\hat{\bm{\alpha}}_{1}\\ \hat{\bm{\alpha}}_{1}^{\dagger}\end{pmatrix}=\hat{\mathcal{U}}_{1}^{\dagger}\begin{pmatrix}\hat{\bm{f}}\\ \hat{\bm{f}}^{\dagger}\end{pmatrix},\quad\begin{pmatrix}\bm{\alpha_{\ell}}\\ \bm{\alpha_{\ell}}^{\dagger}\end{pmatrix}=\hat{\mathcal{U}}_{\ell}^{\dagger}\begin{pmatrix}\hat{\bm{f}}\\ \hat{\bm{f}}^{\dagger}\end{pmatrix},$ (S72)
@@
Given the matrix $\hat{U}$ of eigenvectors of $\hat{H}_{\rm BdG}$, the matrix $\hat{U}^{\prime}$ of eigenvectors of $\hat{H}_{\rm BdG}^{\prime}$ is then $\hat{U}^{\prime}=\hat{Q}\hat{U}$. The latter can then be used to construct the representation of the BCS vacuum |$\ket{\Phi^{\prime}}$ with respect to the reference state |$\ket{\Phi_{1}}$, according to Eqs. (S63) and (III), and then be used to compute its overlap with |$\ket{\Phi}$ via Eq. (S74). However, there is an additional caveat. Using the representation (S63) fixes the gauge of |$\ket{\Phi^{\prime}}$ by imposing a real $\langle{\Phi^{\prime}|\Phi_{1}}\rangle$, which is unwanted. Indeed, consistency with the gauge choice made along the path, $\langle{\Phi_{i}|\Phi_{1}}\rangle\in\mathbb{R}$, only requests that each ground state |$\ket{\Phi}$ (and not also $\hat{Q}\ket{\Phi}$) has real overlap with the reference state |$\ket{\Phi_{1}}$. To remedy for this, in the computation of the phase in Eq. (S62) we introduce a fictitious intermediate step to |$\ket{\Phi_{i}^{\prime}}$ to fix the gauge choice, namely

### v3

@@ II.2 Numerical determination of the Floquet Hamiltonian
As discussed in the main text and detailed in Sec. III, the calculation of the exchange phases builds on the assumption that ground states of the effective Floquet Hamiltonian for the driven fermions, corresponding to a given distribution of $e$ quasiparticles, are prepared. In the high-frequency regime considered here, all $\hat{H}_{F,\mathrm{BdG}}^{[t_{0}]}$ for any initial time $t_{0}$ coincide to leading order in $1/\omega$ [see Eq. (S31b)]: the Hamiltonians are well approximated by the time-averaged Hamiltonian of Eq. (S36), which possesses entirely the desired topological properties for the effective fermion bands, and only differ by small corrections due to the weak higher-order terms. This is indeed verified by comparing the vortex exchange phases obtained using ground states of $\hat{H}_{\rm F,BdG}^{[t_{0}]}$ for different initial time $t_{0}$, as shown in Fig. S4. From Fig. S4, one can appreciate that, for all system sizes shown, the variation in the exchange phase is within $\sim 0.3\%$.
Figure S4: Vortex exchange phase $R_{\psi}^{ee}$ in the $\psi$ sector, divided by its ideal value $83\pi/$ 8, as a function of the global phase $\varphi_{0}$ of the driving functions, defined by $d_{x}(t)=J+2\Delta\cos(\omega t+\varphi_{0})$ and $d_{y}(t)=J+2\Delta\cos(\omega t+\varphi_{0}+\pi/2)$. The relative phase $\pi/2$ imprints $p+ip$ pairing, as discussed in the main text. The driving amplitude and frequency are set at $J=\Delta=\mu_{\psi}/16$ and $\omega=2\mu_{\psi}-4J=7\mu_{\psi}/4$ in the topological region. The gray dotted line is a guide for the eye at the ideal value 1.
@@ III Vortex exchange phases
with a Thouless matrix $T_{ij}$, with respect to the chosen reference state |$\ket{0_{\rm tc}}$, which is the fermion vacuum. The normalization constant is $\mathcal{N}={\mathrm{det}(u)}$, with $u$ defined in Eq. (S47). In case |$\ket{\Phi}$ has vanishing overlap with the fermion vacuum, $\langle\Phi|0_{\rm tc}\rangle=0$, a different reference state needs be chosen, as further discussed below. By imposing $\hat{\alpha}_{n}\ket{\Phi}=0$, the Thouless matrix $T_{ij}$ is found as $T=-(u^{*})^{-1}v^{*}$ [76]. Since, for the toric-code model on a torus, only pairs of fermionic quasiparticles can be created or annihilated, the physical states must have even fermionic parity. The ground state parity is determined by the Floquet-BdG Hamiltonian and is computed following Ref. [4]. Namely, the Floquet Hamiltonian is first expressed in terms of Majorana operators [Eq. (S14)] as $\hat{H}_{F}=i\sum_{p,p^{\prime}}{H}^{(\gamma)}_{F,pp^{\prime}}\hat{\gamma}_{p}\hat{\gamma}_{p^{\prime}}^{\prime}$. The ground state parity is then given by the Pfaffian $\mathrm{Pf}(B)$ of the real skew-symmetric matrix $B$ defined as
@@
In case the lowest energy state of $\hat{H}_{\rm BdG}$ has odd parity, the physical ground state will be the one containing the first BdG quasiparticle. This can be determined as the state which is annihilated by all the operators $\alpha_{j}$ with $j\neq 1$ and by $\alpha_{1}^{\dagger}$ [76, 31].
Geometric phase. To compute the geometric phase accumulated by moving a vortex along a given path, we follow the strategy of Refs. [31, 63], adapted here to the Floquet system. We consider controlled vortex movement within the bulk, that never spans the whole lattice, such that no flipping of the eigenvalues of the Wilson loop operators (change of superselection sector) can occur. For each position of the vortices along the path, the fermion Floquet-BCS corresponding to that vortex configuration is computed from the BdG Hamiltonian as explained in the previous section (see also Sec. II.2), resulting in a sequence of fermion ground states |$\ket{\Phi_{j}}$. To avoid dependence of the phase from the details of the exchange path, Levin-Wen’s protocol [63] is employed, which requires one to compute the difference $\theta=\theta_{P}-\theta_{P^{\prime}}$ of the phases $\theta_{P}$ accumulated along two ‘T-shaped’ paths depicted in Fig. 3(a). The phase along path $P$ is computed as
@@
On the one hand, this choice of reference state ‘saves’ one computation step for $\theta_{P}$ by fixing the phase in the first step to $\mathrm{Arg}\langle\Phi_{2}|\Phi_{1}\rangle=0$, since the ansatz of Eq. (S55) implicitly fixes the gauge along the path through the condition that $\langle\Phi_{i}|\Phi_{1}\rangle=\mathcal{N}_{1}$ is real. On the other, it is a convenient choice since $\langle\Phi_{i}|\Phi_{1}\rangle$ typically remains fairly large, avoiding numerical issues in the construction of the state associated with small overlaps.
@@
$\begin{pmatrix}\hat{\bm{\alpha}}_{1}\\ \hat{\bm{\alpha}}_{1}^{\dagger}\end{pmatrix}=\hat{\mathcal{U}}_{1}^{\dagger}\begin{pmatrix}\hat{\bm{f}}\\ \hat{\bm{f}}^{\dagger}\end{pmatrix},\quad\begin{pmatrix}\hat{\bm{\alpha}}_{\ell}\\ \hat{\bm{\alpha}}_{\ell}^{\dagger}\end{pmatrix}=\hat{\mathcal{U}}_{\ell}^{\dagger}\begin{pmatrix}\hat{\bm{f}}\\ \hat{\bm{f}}^{\dagger}\end{pmatrix},$ (S56)
@@
Given the matrix $\hat{U}$ of eigenvectors of $\hat{H}_{\rm BdG}$, the matrix $\hat{U}^{\prime}$ of eigenvectors of $\hat{H}_{\rm BdG}^{\prime}$ is then $\hat{U}^{\prime}=\hat{Q}\hat{U}$. The latter can then be used to construct the representation of the BCS vacuum |$\ket{\Phi^{\prime}}$ with respect to the reference state |$\ket{\Phi_{1}}$, according to Eqs. (S55) and (S57), and then be used to compute its overlap with |$\ket{\Phi}$ via Eq. (S58). However, there is an additional caveat. Using the representation (S55) fixes the gauge of |$\ket{\Phi^{\prime}}$ by imposing a real $\langle{\Phi^{\prime}|\Phi_{1}}\rangle$, which is unwanted. Indeed, consistency with the gauge choice made along the path, $\langle{\Phi_{i}|\Phi_{1}}\rangle\in\mathbb{R}$, only requests that each ground state |$\ket{\Phi}$ (and not also $\hat{Q}\ket{\Phi}$) has real overlap with the reference state |$\ket{\Phi_{1}}$. To remedy for this, in the computation of the phase in Eq. (S54) we introduce a fictitious intermediate step to |$\ket{\Phi_{i}^{\prime}}$ to fix the gauge choice, namely
@@
IV Driving-induced heating
Periodically-driven quantum systems are generically expected to heat up to a featureless infinite-temperature-like state at long times as a consequence of energy absorption from the drive [45, 78, 46]. In this section, we analyze these aspects for the driven model studied in this work, elaborating on the statements given in the main text. First of all, since the local state space (spins) is bounded, one expects that the timescale over which heating kicks in can be prolonged exponentially with increasing driving frequency [43, 44]. Moreover, a key observation is that our driven toric code maps to a model of static $e$ boson and driven noninteracting fermions in the quasiparticle picture. Such integrable systems are predicted not to heat up to a featureless state, due to the existence of an extensive number of conserved quantities given by the occupations of the single-particle Floquet modes [47, 48], although the expectation value of certain observables may still resemble the infinite-temperature case [79]. In the following, we present numerical results confirming that no heating takes place in the high-frequency regime of interest, whereas a strong energy increase is observed away from it when vortices are present.
IV.1 Numerical analysis
To investigate the deviation of the exact Floquet dynamics from the time averaged description, resulting in heating, we will inspect heating measures for varying ratio $J/\omega$ between the driving amplitudes $J=\Delta$ and the driving frequency $\omega$, which is the relevant perturbative parameter in the high-frequency expansion. Recall that, in the time-averaged model of Eq. (5), the detuning of the driving frequency $\omega$ from the fermion-pair gap $2\mu_{\psi}$ determines the effective chemical potential $\mu=\omega/2-\mu_{\psi}$ for the driven fermions. We analyse a regime in which the time-averaged model is in the topological phase, namely $\mu=-2J$, implying $\omega=2\mu_{\psi}-4J$. Thus, to maintain non-vanishing $\omega$, we will consider the range $J/\omega\leq 1$.
To investigate the impact of heating during the time evolution, we first consider the stroboscopic dynamics |$\ket{\psi(nT)}$ with the system initially prepared in the fermion ground state |$\ket{\Psi_{0}}$ of the time-averaged Hamiltonian $\hat{H}_{\rm avg}$ of Eq. (5), for fixed background distributions of $e$ particles. The fact that the exact Floquet Hamiltonian differs from the time-averaged one will lead to deviations in the expectation value of physical observables, which may be considered as heating [37]. To verify this, we monitor heating through the quantity [80]
$\mathcal{Q}(nT)=\frac{\bra{\psi(nT)}\hat{H}_{\rm avg}\ket{\psi(nT)}-E_{0}}{E_{\infty}-E_{0}},$ (S62)
where $E_{0}=\bra{\Psi_{0}}\hat{H}_{\rm avg}\ket{\Psi_{0}}$ is the initial and ideal value, and $E_{\infty}=(1/2^{N})\mathrm{tr}(\hat{H}_{\rm avg})$ is the infinite-temperature value. Thus, $\mathcal{Q}(nT)=0$ if no heating occurs, whereas $\mathcal{Q}(nT)=1$ if an infinite-temperature-like value is reached. Technical details about the calculation of $\mathcal{Q}(nT)$ are given in Sec. IV.2 in the following. The results for $L=24$ are shown in Fig. S5, for different ratio $J/\omega$. The panel S5(a) depicts the result in the absence of vortices, whereas S5(b) corresponds to the average of 50 random configurations of a single pair of $e$ particles. Recall that the presence of vortices impacts the dynamics by inducing anyonic effects as the fermions move. For ratios in in the high-frequency regime ($J/\omega\ll 1$), $\mathcal{Q}(nT)$ saturates well below the infinite-temperature value both in the presence and in the absence of vortex pairs with similar behaviour. Moreover, as predicted, heating is strongly suppressed with decreasing $J/\omega$. The time-averaged Hamiltonian thus remains a good description of the system even at long times. A different behaviour is observed when $J/\omega\sim 1$, corresponding to the regime where the high-frequency expansion starts to diverge: the value of $\mathcal{Q}(nT)$ is much larger instead and approaches the infinite-temperature value.
Figure S5: Stroboscopic value of $\mathcal{Q}(nT)$ of Eq. (S62) as a function of the number of Floquet period, in the absence (a) and presence (b) of a vortex pair, for antiperiodic boundary conditions for the fermions. The results in (b) are averaged over 50 random $e$-pair configurations. The standard deviation obtained is barely visibile on the scale of the plot, indicating little dependence on the vortex position, and we thus did not include it. The thin gray line is a guide for the eye at value 1.
To investigate the long-time dynamics more systematically, we analyze the infinite-time average [79]
$\overline{\mathcal{Q}}=\frac{\bra{\Psi_{0}}\overline{H}_{\rm avg}\ket{\Psi_{0}}-E_{0}}{E_{\infty}-E_{0}}.$ (S63)
The Hamiltonian $\overline{H}_{\mathrm{avg}}$ is obtained from $\hat{H}_{\rm avg}$ by neglecting the off-diagonal matrix elements in the basis of Floquet modes. This is reported in Fig. S6 as a function of the ratio $J/\omega$ for different system size $L=8,16,24$. The panel S6(a) reports the results in the absence of vortices. In S6(b), we investigate the impact of vortices by considering random distributions of $e$ particles at a given density $\rho_{v}$ and then averaging $\overline{Q}$ over 20 configurations for each $\rho_{v}$ and $L=16$. In the high-frequency regime, $\overline{\mathcal{Q}}$ is always close to zero independently from the presence of vortices: the system does not heat up, confirming the practicality of the Floquet engineering scheme proposed. In the anomalous Floquet regime, where the time-averaged description is not valid anymore, $\overline{\mathcal{Q}}$ increases significantly, but still saturates below the infinite-temperature state in the absence of vortices and exhibits little dependence on the system size. The situation is different in the presence of vortices, where values closer to $\overline{\mathcal{Q}}=1$ are found. This indicates that, although the many-body occupation of the single-particle Floquet modes is conserved, the dynamics with vortices strongly mixes the single-particle eigenstates of $\hat{H}_{\rm avg}$, yielding infinite-temperature-like expectation values of the one-body observable $\hat{H}_{\rm avg}$ at long times [79].
Figure S6: Infinite-time average $\overline{\mathcal{Q}}$ of Eq. (S63) as a function of the ratio $J/\omega$ for different system size in the absence of vortices (a) and for $L=16$ and different vortex densities $\rho_{v}$ (b), with antiperiodic fermion boundary conditions. The results in (b) are averaged over 20 random vortex configurations at the given density (b). The standard deviation obtained is barely visibile on the scale of the plot, indicating little dependence on the vortex position, and we thus did not display it.
IV.2 Calculation of the heating measures
In this subsection, we explain how the figures of merit $\mathcal{Q}(nT)$ and $\overline{\mathcal{Q}}$ of Eqs. (S62) and (S63) used to quantify heating in Sec. IV.1 are computed. The key quantities needed are the expectation values
$\mathrm{(i)}\bra{\Psi_{0}}\hat{U}(nT)^{\dagger}\hat{H}_{\rm avg}\hat{U}(nT)\ket{\Psi_{0}},$ (S64a)
$\mathrm{(ii)}\bra{\Psi_{0}}\overline{{H}}_{\rm avg}\ket{\Psi_{0}}.$ (S64b)
which are of the general form $\bra{\Phi}\hat{H}\ket{\Phi}$ for a Hamiltonian $\hat{H}$ and a BCS state |$\ket{\Phi}$. To compute such quantities, consider the BdG representation of $\hat{H}$, $\hat{H}=(1/2)(\hat{\bm{f}}^{\dagger},\hat{\bm{f}})\hat{H}^{\rm BdG}(\hat{\bm{f}},\hat{\bm{f}}^{\dagger})^{T}$ and the quasiparticle operators $(\hat{\bm{\alpha}},\hat{\bm{\alpha}}^{\dagger})^{T}=\hat{\mathcal{U}}(\hat{\bm{f}},\hat{\bm{f}}^{\dagger})^{T}$ which annihilate the BCS state |$\ket{\Phi}$ (see Sec. III), where $\hat{\mathcal{U}}$ has the form of Eq. (S47), namely
$\hat{\mathcal{U}}=\begin{pmatrix}u&v^{*}\\ v&u^{*}\end{pmatrix}.$ (S65)
The expectation value $\langle\hat{H}\rangle=\bra{\Phi}\hat{H}\ket{\Phi}$ then reads as
$\langle\hat{H}\rangle=\frac{1}{2}\sum_{i,j=1}^{N}\hat{H}^{\rm BdG}_{i,j}\langle\hat{f}_{i}^{\dagger}\hat{f}_{j}\rangle+\frac{1}{2}\sum_{i,j=1}^{N}\hat{H}^{\rm BdG}_{i,j+N}\langle\hat{f}_{i}^{\dagger}\hat{f}_{j}^{\dagger}\rangle+\frac{1}{2}\sum_{i,j=1}^{N}\hat{H}^{\rm BdG}_{i+N,j}\langle\hat{f}_{i}\hat{f}_{j}\rangle+\frac{1}{2}\sum_{i,j=1}^{N}\hat{H}^{\rm BdG}_{i+N,j+N}\langle\hat{f}_{i}\hat{f}_{j}^{\dagger}\rangle.$ (S66)
Expressing the operators $(\hat{\bm{f}}^{\dagger},\hat{\bm{f}})$ in terms of $(\hat{\bm{\alpha}}^{\dagger},\hat{\bm{\alpha}})$ by means of $\hat{\mathcal{U}}$ and Eq. (S65), one then finds
$\langle\hat{f}_{i}^{\dagger}\hat{f}_{j}\rangle=(vv^{\dagger})_{ij},$ $\langle\hat{f}_{i}^{\dagger}\hat{f}_{j}^{\dagger}\rangle=(uv^{\dagger})_{ij},$
$\langle\hat{f}_{i}\hat{f}_{j}\rangle=(vu^{\dagger})_{ij},$ $\langle\hat{f}_{i}\hat{f}_{j}^{\dagger}\rangle=(uu^{\dagger})_{ij}\ .$ (S67)
Returning to the specific expectation values in Eqs. (S64a) and (S64b), the Hamiltonian $\overline{H}_{\rm avg}$ of (S64a) is obtained, as discussed in Sec. IV.1, by dropping the off-diagonal elements of $\hat{H}_{\rm avg}$ in the basis of Floquet states. The Hamiltonian $\hat{U}(nT)^{\dagger}\hat{H}_{\rm avg}\hat{U}(nT)$ of (S64a) is computed from the numerically-determined end-of-period propagator $\hat{U}(T)$ and Floquet theorem, implying $\hat{U}(nT)=[\hat{U}(T)]^{n}$.